# A Generation Framework with Strict Constraints for Crystal Materials Design

### v2

@@ 1 Introduction
Designing novel and high-performance crystal structures has long been recognized as a critical research direction in materials science. Traditional crystal materials design relies on computational simulations, such as Density Functional Theory (DFT) (1) and high-throughput calculations, which demand substantial human expertise and computational resources (2). Over the past few decades, the application of computer technology in the field of crystal structure generation has made systematic new materials design a reality (3), leading to the emergence of a series of representative methods, such as simulated annealing (4), random search (5), genetic algorithms (6), and particle swarm optimization (7). Nevertheless, these methods still exhibit limitations in generating valid and diverse crystal structures.
In recent years, large open-source material databases (8; 9; 10) have greatly advanced data-driven methods in crystal generation (11; 12). Among these, generative models have been proved particularly effective in generating valid and diverse crystal structures (13; 14). However, constraint-based targeted crystal structure generation has remained challenging. CrystalFormer (15) generates space group-controlled crystal materials using space groups as initial constraint, while DiffCSP++ (16) produces strictly symmetry-compliant crystals based on elemental and symmetry constraints. MatterGen (17) proposes a conditional diffusion model capable of guiding materials generation with specified constraints, including properties and chemical composition, though these constraints remain flexible and do not always fully meet the preset conditions.
@@
2 Related works
Crystal Structure Generation
Crystal structure generation methods based on data-driven have witnessed rapid advancements. CDVAE (13) combines the variational autoencoder (VAE) with diffusion model, first generating atomic types and lattice vectors from encoded crystal features, and then using these as conditional inputs to the diffusion model to generate new crystal structures. Building upon CDVAE, Con-CDVAE (18) incorporates material properties as conditions for crystal generation and designs a two-step training method to align the encoded features of properties. DiffCSP (14) optimizes the generation of lattice matrices and atomic coordinates through a joint diffusion framework, while its successor DiffCSP++ (16) introduces constraints based on crystal symmetry, using symmetry basis matrices to constrain the lattice vectors and Wyckoff position coordinates to constrain the atomic coordinates, ensuring that the generated crystal structures strictly adhere to the input crystal symmetry. MatterGen (17) adopts a joint diffusion framework, uses polar decomposition to represent the lattice as O(3)-invariant symmetric matrices, and further introduces an adapter module for fine-tuning based on specific property constraints. UniMat (19) proposes a unified method to represent any crystal structure to address the compression problem of crystal information, and trains a diffusion model on this representation. Compared to the aforementioned methods, our proposed model generates more desirable and accurate crystal structures by incorporating both input and self-generated constraints.
Crystal Symmetry Generation
Crystal symmetry, encompassing space groups and Wyckoff positions, plays a pivotal role in the description of crystal structures. We demonstrate the effects of symmetry on the lattice geometry and atomic arrangements in Table 5 and Figure 6 respectively. Determining symmetry simplifies the complexity of generating crystal structures. Zhao et al. (20) proposed a neural network model that combined Random Forest (RF) and Multi-Layer Perceptron (MLP), and achieved effective prediction of crystal systems and space groups by incorporating features such as Magpie, atomic vectors, and one-hot encoding. Wang et al. (21; 22) proposed a space group prediction method based on crystal chemical formulas and crystal-related feature descriptors (such as total atomic count, average atomic radius, maximum sound velocity, minimum polarizability, and other properties). Rewc-GNN (23) aims to predict the space groups of porphyrin crystals by encoding the atomic properties and interatomic interactions within the crystal through the integration of the Rewc algorithm and Graph Neural Networks. Learning-based crystal space group prediction models have demonstrated significant potential in prediction accuracy and generalization capabilities, providing robust support for the discovery and design of novel crystal structures. However, based on our current literature review, research on LLM-based symmetry prediction, and studies focusing on Wyckoff position prediction within crystal symmetry, remains relatively scarce.
Large Language Models
Over the past few years, LLMs have undergone rapid iteration and evolution (24; 25), gradually approaching and even surpassing human-level performance across various complex benchmarks. Alongside proprietary models, open-source counterparts have achieved remarkable progress, including the DeepSeek series  (26; 27), LLaMA series (28; 29), Qwen series (30; 31), and GLM series (32; 33). These open-source models are progressively narrowing the performance gap with their closed-source counterparts. LLMs have also found broad applications in materials science: MatChat (34)generates high-confidence solid-state reaction pathways by analyzing inorganic materials synthesis data from tens of thousands of research papers. The MatterChat (35)model achieves breakthrough capabilities by integrating atomic structural data with natural language processing, bridging machine-learned interatomic potentials with pretrained LLMs to enable full-resolution atomic structure analysis. The LLMatDesign (36)framework leverages LLM agents for zero-shot prediction of material properties and novel materials design. The GPTFF (37)universal pretrained force field, incorporating Transformer attention mechanisms, delivers precise predictions of interatomic interactions. CrystalLLM (38) employs LLMs to generate crystal structures. However, since learning and generating extensive atomic coordinates fall outside the core competencies of LLMs, this paper proposes the development of a crystal constraint generator powered by LLMs.
@@ 2 Related works
3 Our method : CrystalGF
In this section, we present CrystalGF, a novel framework for crystal materials generation that synthesizes desired crystal structures guided by chemical composition and material properties.
3.1 Architecture
A crystal generation model conditioned with chemical composition and material properties could be described as ${M}=H(C,P)$, where $H(.)$ represents a crystal structure generator, $C=({c_{1}},{c_{2}},...,{c_{k}})$ are the chemical composition constraint in which k components are specified, P is property constraint. $M=(A,X,L)$ represents for crystal structure, where $A=[{a_{1}},{a_{2}},...,{a_{n}}]$ are the elemental composition of atoms inside a unit cell, $X=[{x_{1}},{x_{2}},...,{x_{n}}]\in{{[0,1)}^{3\times n}}$ are fractional coordinates of $n$ atoms, $L=[{l_{1}},{l_{2}},{l_{2}}]\in{\mathbb{R}^{3\times 3}}$ is the lattice. In this work, we formulate our model as:
$\begin{array}[]{c}{M}=H(G(C,P),P)\end{array}$
Notice that for our model, the actual input constraints are still chemical composition and material properties, and we utilize a constraint generator to obtain additional constraints. We utilize a LLMs-based model as our constraint generation module (Detailed in §3.2), and a multi-conditional constrained diffusion model as the crystal structure generation module (Detailed in §3.3). Furthermore, both H(.) and G(.) could be any model as long as the inputs and outputs match.
3.2 Constraint Generator
We design the following three tricks in the prompts to regulate and enhance the generation results. (1) Generation scope constraints: For space groups, we explicitly specify in the "Instruction" part that valid space group types should range from 0 to 230; For Wyckoff positions, we provide candidate Wyckoff letters in the "Input" part. (2) Numerical standardization: To minimize the impact of text length on LLMs, we standardized all input material properties by retaining three decimal places and adaptively determining appropriate units based on property types. (3) Text optimization: Instead of directly generating target space group type and Wyckoff positions, we incorporate the phrase "the most suitable…" into the prompts, which has been empirically shown to improve generation accuracy.
3.3 Crystal Structure Generator
The crystal structure generator is designed to receive symmetry constraints, elemental composition, and target properties generated by constraint generator, subsequently generating desired crystal structures. We employ a diffusion model to jointly generate the lattice matrix $\bm{L}$ and fractional coordinates $\bm{F}$ under these conditions. Detailed descriptions of the forward diffusion process and reverse generation procedure are provided in the B.1.
The overall architecture of our denoising model is illustrated in Figure 3, which derives two denoising terms $\hat{\epsilon}_{k}$ , $\hat{\epsilon}_{{F}^{{}^{\prime}}}$ under the constraints of elemental composition, crystal symmetry, and target properties. For feature embedding, we introduce continuous material property embeddings into input feature representations. Continuous properties, following Con-CDVAE (18), are expanded by a Gaussian Radial Basis Function (RBF) and embedded by an MLP:
For the embedding of noised lattices and noised coordinates, we adopt the approach described in DiffCSP++ (16). Subsequently, we employ a multi-level cascade decoder to generate denoising terms. To enhance the model’s fitting capacity, we incorporate a multi-head cross-attention mechanism within the decoder architecture to fuse information from both the noised feature and the input feature. The structure of the decoder in the n-th layer is designed as
@@ 3.3 Crystal Structure Generator
4 Experiments
We conducted three sets of experiments to validate the effectiveness of our framework. First, to verify the impact of the constraint generator on final crystal structure generation, we selected four different generators and demonstrated a negative correlation between symmetry accuracy and property error (Detailed in §4.1). Second, we validated the effectiveness of our multi-conditional crystal structure generator (Detailed in § 4.2). Finally, to confirm our framework’s capability to accurately generate crystal materials satisfying target property requirements, we conducted comparative experiments with MatterGen and Con-CDVAE (Detailed in §4.3).
@@ 4.1 Crystal Constraints Generation
We picked 34507 samples from the Materials Project (MP) (8) database, all of which were stable crystal structures, whose energy per atom after relaxation via DFT is within 0.1 eV per atom above the convex hull, with less than 20 atoms in crystal unit cell. Each sample includes a crystal structure and its associated properties, such as band gap, formation energy, energy above hull.
@@
The distribution of crystal space group types in our dataset is visualized in Figure 4. Given the extreme sparsity for certain space group types, for instance, only 15 samples for $I2_{1}2_{1}2_{1}$ (No.24) and 0 sample for $Ia\text{-}3d$ (No.230), and the broad range of space group types, we adopted a strategy inspired by previous researchers (21; 22) to facilitate statistical analysis and presentation of model inference results: during model training, all original labels are preserved, but during evaluation, space group types with fewer than 500 instances are aggregated into an “other” category for statistical clarity. This approach is extended to Wyckoff positions generation, where Wyckoff letters with fewer than 50 occurrences are similarly categorized as “other”.
We selected four widely downloaded open-source LLMs under 10B as the base models for the generator and fine-tuned them using the Llama-Factory (39) framework. We provided more details in Appendix B.1. The results of the crystal symmetry generation are presented in Table 1. In experiments using elemental composition and material property as input, the highest generation accuracies for space groups and Wyckoff positions reached 35.15% and 62.58% respectively, both obtained using the Llama-based generator. When using chemical formula and material property as input, the peak accuracies for space groups and Wyckoff positions improved to 66.73% and 82.83%, achieved by the Llama-based and GLM-based generators respectively. Correspondingly, when employing four constraint generators based on different LLMs and specifying bandgap and formation energy as property constraints, the error probability distributions between the simulated properties of crystals generated by CrystalGF and target properties are presented in Tables 3 and Table 4. Combined with the results from Table 1, it demonstrates that higher symmetry accuracy achieved by constraint generators corresponds to smaller deviations between the generated crystal structures and target properties.
@@ 4.2 Crystal Structure Generation
We evaluated our crystal structure generation module on three publicly available datasets with distinct data distributions: Perov-5, MP-20, and MPTS-52. Perov-5  (40) comprises 18,928 perovskite crystals exhibiting structural similarity but compositional diversity, each containing 5 atoms per unit cell. MP-20  (41) contains 45,231 samples representing experimentally realized crystals with $\leq$ 20 atoms per unit cell. MPTS-52 serves as a more challenging extension of MP-20, consisting of 40,476 structures with up to 52 atoms per unit cell. For all datasets, we adopt the data splitting protocols established in prior works  (13; 14) to ensure benchmarking consistency. The hyper-parameters and training details are summarized in Appendix B.1.
For each generated crystal structure, we calculate the match rate through the Structure Matcher class in Pymatgen (42), with thresholds stol = 0.5, angle tol = 10, ltol = 0.3, following previous work. The match rate represents the ratio of matched structures relative to the total number within the testing set, and the RMSE is averaged over the matched pairs, and normalized by $3]{V/N}$, where $V$ is the volume of the lattice, $N$ is the atom number of crystals. We compared our approach with the following two generation-based methods, including CDVAE (13) and DiffCSP++ (16). The results of CDVAE and DiffCSP++ are obtained by reproducing the inference process using publicly released datasets and pretrained model weights.
@@ 4.3 Material Properties Comparison
In this subsection, we compared the differences between the simulated properties of the generated structures and the target properties. We employ density functional theory (DFT) for computational simulations using the Vienna Ab initio Simulation Package (43) (44) (VASP), more details in Appendix B.3. For comparative analysis, we fine-tuned and retrained two multi-conditional crystal generation models: MatterGen and Con-CDVAE, the training details are provided in the appendix B.1. In practical crystal materials design, it is considered acceptable if the absolute error of the properties is less than 0.05, and meet the requirement if less than 0.01, and the unit of band gap and formation energy are eV and eV/atom, respectively. Therefore, we compared the probability distributions of property absolute errors between different models in Tables 3 and Table 4.
@@
5 Conclusion
In this work, we propose CrystalGF, a generation framework with strict constraints for crystal generation. We decompose the generation process into intermediate constraint generation and structural generation. Within our framework, intermediate constraints such as symmetry and component ratios are generated by a LLMs-based constraint generator according to chemical composition and target properties, which are strictly enforced by a diffusion-based crystal generator to ensure constraint-controlled structure generation. Experimental results demonstrate that compared with previous methods, structures produced by our method are more than twice as likely to meeting the target properties, while maintaining rigorous adherence to input chemical composition. We believe that the multiple strict constraint ability of our method represent an important advance towards automaticly and accurately designing of crystal materials. Furthermore, it could could be extended in many directions, such as covering more physical and chemical properties or transfering to other reaearch areas like molecules generation.
@@ 5 Conclusion
Appendix A Crystal Supplementary Know

### v3

@@ Abstract
keywords:
crystal materials, multiple constraints, target properties
@@ 1 Introduction
Designing novel and high-performance crystal structures has long been recognized as a critical research direction in materials science. Traditional crystal materials design relies on computational simulations, such as Density Functional Theory (DFT) [orio2009density] and high-throughput calculations, which demand substantial human expertise and computational resources [yao2023machine]. Over the past few decades, the application of computer technology in the field of crystal structure generation has made systematic new materials design a reality [graser2018machine], leading to the emergence of a series of representative methods, such as simulated annealing [timmermann2020iro], random search [yang2021exploration], genetic algorithms [cheon2020crystal], and particle swarm optimization [2022CSPTHAC]. Nevertheless, these methods still exhibit limitations in generating valid and diverse crystal structures.
In recent years, large open-source material databases [jain2013commentary, schmidt2022large, schmidt2022dataset] have greatly advanced data-driven methods in crystal generation [nouira2018crystalgan, ren2022invertible]. Among these, generative models have been proved particularly effective in generating valid and diverse crystal structures [xie2021crystal, jiao2024crystal, miller2024flowmm, gruver2024fine, wu2025periodic]. However, constraint-based targeted crystal structure generation has remained challenging. CrystalFormer [cao2024space] generates space group-controlled crystal materials using space groups as initial constraint, while DiffCSP++ [jiao2024space] produces strictly symmetry-compliant crystals based on elemental and symmetry constraints. MatterGen [zeni2023mattergen] proposes a conditional diffusion model capable of guiding materials generation with specified constraints, including properties and chemical composition, though these constraints remain flexible and do not always fully meet the preset conditions.
@@
2 Results
2.1 Crystal Generation Framework
In CrystalGF, we design a two-stage framework for generating crystal materials (Figure 1). Many performance-oriented crystal generation methods overlook the influence of internal structural characteristics on material properties, whereas our approach establishes a connection between macroscopic performance and microstructure through a constraint generator (Appendix B.2) and a structure generator (Appendix B.3). Specifically, the constraint generator learns the relationship between the macroscopic properties and internal structural features of crystals to generate corresponding structural constraints, including space groups and Wyckoff positions of atoms. The structure generator, guided by multiscale constraints such as macroscopic properties and internal structural characteristics, produces crystal structures that meet the target performance criteria.
We conducted three sets of experiments to validate the effectiveness of our framework. First, to verify the impact of the constraint generator on final crystal structure generation, we selected four different generators and demonstrated a negative correlation between symmetry accuracy and property error (Section 2.2). Second, we validated the effectiveness of our multi-conditional crystal structure generator (Section 2.3). Finally, to confirm our framework’s capability to accurately generate crystal materials satisfying target property requirements, we conducted comparative experiments with MatterGen and Con-CDVAE (Section 2.4).
@@ 2.2 Crystal Constraints Generation
We picked 34507 samples from the Materials Project (MP) [jain2013commentary] database, all of which were stable crystal structures, whose energy per atom after relaxation via DFT is within 0.1 eV per atom above the convex hull, with less than 20 atoms in crystal unit cell. Each sample includes a crystal structure and its associated properties, such as band gap, formation energy, energy above hull.
@@
The distribution of crystal space group types in our dataset is visualized in Figure 2. Given the extreme sparsity for certain space group types, for instance, only 15 samples for $I2_{1}2_{1}2_{1}$ (No.24) and 0 sample for $Ia\text{-}3d$ (No.230), and the broad range of space group types, we adopted a strategy inspired by previous researchers [che2023crystal, wang2023crystallographic] to facilitate statistical analysis and presentation of model inference results: during model training, all original labels are preserved, but during evaluation, space group types with fewer than 500 instances are aggregated into an “other” category for statistical clarity. This approach is extended to Wyckoff positions generation, where Wyckoff letters with fewer than 50 occurrences are similarly categorized as “other”.
We selected four widely downloaded open-source LLMs under 10B as the base models for the generator and fine-tuned them using the Llama-Factory [zheng2024llamafactory] framework. We provided more details in Appendix C.1. The results of the crystal symmetry generation are presented in Table 1. In experiments using elemental composition and material property as input, the highest generation accuracies for space groups and Wyckoff positions reached 35.15% and 62.58% respectively, both obtained using the Llama-based generator. When using chemical formula and material property as input, the peak accuracies for space groups and Wyckoff positions improved to 66.73% and 82.83%, achieved by the Llama-based and GLM-based generators respectively. Correspondingly, when employing four constraint generators based on different LLMs and specifying bandgap and formation energy as property constraints, the error probability distributions between the simulated properties of crystals generated by CrystalGF and target properties are presented in Tables 5 and Table 6. Combined with the results from Table 1, it demonstrates that higher symmetry accuracy achieved by constraint generators corresponds to smaller deviations between the generated crystal structures and target properties.
@@ 2.3 Crystal Structure Generation
We evaluated our crystal structure generation module on three publicly available datasets with distinct data distributions: Perov-5, MP-20, and MPTS-52. Perov-5  [castelli2012new] comprises 18,928 perovskite crystals exhibiting structural similarity but compositional diversity, each containing 5 atoms per unit cell. MP-20  [jain2013materials] contains 45,231 samples representing experimentally realized crystals with $\leq$ 20 atoms per unit cell. MPTS-52 serves as a more challenging extension of MP-20, consisting of 40,476 structures with up to 52 atoms per unit cell. For all datasets, we adopt the data splitting protocols established in prior works  [xie2021crystal, jiao2024crystal] to ensure benchmarking consistency. The hyper-parameters and training details are summarized in Appendix C.1.
For each generated crystal structure, we calculate the match rate through the Structure Matcher class in Pymatgen [ong2013python], with thresholds stol = 0.5, angle tol = 10, ltol = 0.3, following previous work. The match rate represents the ratio of matched structures relative to the total number within the testing set, and the RMSE is averaged over the matched pairs, and normalized by $\sqrt[3]{V/N}$, where $V$ is the volume of the lattice, $N$ is the atom number of crystals. We compared our approach with the following two generation-based methods, including CDVAE [xie2021crystal] and DiffCSP++ [jiao2024space]. The results of CDVAE and DiffCSP++ are obtained by reproducing the inference process using publicly released datasets and pretrained model weights.
@@
2.4 Structures Metrics Calculation
We employ the structural metrics established by CDVAE [xie2021crystal] and the S.U.N. evaluation tool provided by MatterGen [zeni2023mattergen] to compare different methods. The evaluation metrics from CDVAE [xie2021crystal] include empirical validity, coverage of the test set, and the Wasserstein distances between the distributions of three properties of the generated samples and the test set: atomic density ($d_{\rho}$), number of unique elements ($d_{elem}$), and predicted formation energy ($d_{E}$). For each method, we generate 10,000 samples for evaluation. The results are presented in Table 3. CrystalGF performs comparably to other methods across most metrics, while exhibiting a higher probability of generating valid crystal structures, which can be attributed to its high-quality constrained priors and precise structure generation.
Validity (%) (↑) Coverage (%) (↑) Property Distribution (↓)
MP-20 Struct. Comp. Recall Precision $d_{\rho}$ $d_{E}$ $d_{elem}$
CDVAE 99.97 85.61 99.31 99.47 0.70 0.24 1.28
DiffCSP++ 99.44 86.50 99.72 99.61 0.12 0.05 0.33
FlowMM 96.67 83.25 99.49 99.58 0.23 0.09 0.08
SymmCD 90.34 85.81 99.58 97.76 0.23 0.21 0.40
Con-CDVAE 99.97 81.77 98.29 99.39 0.67 0.15 0.73
MatterGen 99.56 84.50 99.50 99.33 0.26 0.17 0.11
CrystalGF 99.85 90.76 99.31 99.65 0.17 0.07 0.09
Table 3: Results for comparing the validity, coverage, and property distribution metrics.
Stable, unique and novel (S.U.N.) is a critical metric for evaluating the efficiency and cost-effectiveness in discovering new materials during model inference. In this study, to reduce the computational cost of structural relaxation, we use the validation tools provided by MatterGen to compute the S.U.N. values. Thermodynamic stability is determined by calculating the convex hull energy of the materials, where the convex hull represents the lowest-energy mixing state of the material. We employ a pre-trained mattersim[yang2024mattersim] model to predict the energy of generated crystals and run structural relaxation to achieve higher stability. Novelty validation is conducted using the Alex-MP-ICSD reference dataset. According to our research, the crystal system has a significant impact on the S.U.N. metrics. Highly symmetric crystal systems, such as hexagonal and cubic, possess abundant symmetry operations, exhibit greater structural stability, and occur more frequently in nature. In contrast, crystal systems with lower symmetry, such as triclinic and monoclinic, have fewer symmetry operations and looser structural constraints, making them more prone to generating novel structures. However, crystals with low-symmetry space groups are relatively scarce in nature and pose greater synthetic challenges. even if such crystals could be designed theoretically, their practical utility remains limited. Therefore, to balance the influence of different crystal systems, we proposed a balanced S.U.N. metrics, the calculation formula is presented below:
$\text{S.U.N.}_{\text{balanced}}=\frac{1}{k}\sum_{c\in S}\text{S.U.N.}_{c}$ (1)
where $k$ denotes the number of crystal systems with generated structures, and $S$ represents the set of all crystal systems. For each crystal system $c\in S$, $\text{S.U.N.}_{c}$ denotes the S.U.N. score of all structures within that crystal system.
As shown in Table 4, CrystalGF outperforms SymmCD and FlowMM in terms of the S.U.N. metric. Due to the strong constraint framework adopted in our approach, the generated crystals adheres to the structural characteristics (e.g., symmetry) of many existing crystals. As a result, compared with methods such as DiffCSP++ and Symmcd, CrystalGF produces structures with relatively lower novelty, while exhibiting a significant advantage in stability. Furthermore, both Con-CDVAE and MatterGen demonstrate substantially superior S.U.N. performance compared to other methods. However, in the balanced S.U.N. metric, our CrystalGF exhibit minor improvements, whereas Con-CDVAE and MatterGen demonstrate more pronounced decreases. Combined with the space group distribution of crystal structures generated by each method, as illustrated in Figure 4, it is evident that a significant proportion of the structures generated by Con-CDVAE and MatterGen belong to space group 1. Although their overall S.U.N. metrics are relatively high, the lack of symmetry constraints limits their ability to discover S.U.N. high-symmetry structures. In contrast, CrystalGF exhibit a more uniform distribution across space groups from 1 to 230, thereby enabling the exploration of novel structures across a broader range of structural characteristics.
Figure 4: Proportions of space groups for structures generated by each method. The width of each colored segment represents the proportion of crystals possessing that symmetry.
Method Unique Novel Relaxed Relaxed Balanced
Stable S.U.N. S.U.N.
CDVAE 99.37% 71.66% 38.43% 14.45% 7.23%
DiffCSP++ 98.57% 32.82% 70.90% 9.55% 11.17%
FlowMM 83.35% 93.43% 9.64% 4.18% 1.74%
SymmCD 97.73% 74.23% 30.34% 8.37% 8.63%
Con-CDVAE 91.15% 76.81% 46.09% 25.03% 12.53%
MatterGen 98.97% 78.21% 46.19% 27.02% 7.58%
CrystalGF 96.87% 39.67% 64.67% 9.38% 11.14%
Table 4: Percent of S.U.N. samples produced from 10,000 generated crystals for each method.
@@ 2.5 Material Properties Comparison
In this subsection, we compared the differences between the simulated properties of the generated structures and the target properties. We employ density functional theory (DFT) for computational simulations using the Vienna Ab initio Simulation Package [kresse1996efficient] [kresse1996efficiency] (VASP), more details in Appendix C.3. For comparative analysis, we fine-tuned and retrained two multi-conditional crystal generation models: MatterGen and Con-CDVAE, the training details are provided in the Appendix C.1. In practical crystal materials design, it is considered acceptable if the absolute error of the properties is less than 0.05, and meet the requirement if less than 0.01, and the unit of band gap and formation energy are eV and eV/atom, respectively. Therefore, we compared the probability distributions of property absolute errors between different models in Tables 5 and Table 6.
@@
Appendix A Crystal Supplementary Knowledge